OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = 5/7x - 12 A. -5x + 7y = -12 B. -5x - 7y = -84 C. -5x + 7y = -84 D. 5x + 7y = -84

OpenStudy (anonymous):

can someone help me?

OpenStudy (rsadhvika):

y = 5/7x - 12 7y = 5x - 84 -5x + 7y = -84

OpenStudy (anonymous):

thank u